A particle oscillates as per the equation x = 7cos(0.5πt) m, the time taken by the particle to move from the mean position to a point 3.5 m away is

Grade:12th pass

2 Answers

Vikas TU
7 years ago
at 3.5 m time = T/8
Thus T = 2pi/w
w = 0.5pi
Time Period = 2pi/0.5pi = > 4s
Time taken  by the particle to move from the mean position to a point 3.5 m away is:
4/8 = 0.5s.
Sabita kumari
4 years ago
x = A cos wt 
T = 2(pi) / w = 2(pi) / 0.5 pi = 4 secs 
T/4 = 4/4 secs = 1 sec 
Hence , the correct answer is 1 second .